OpenStudy (anonymous):

3√2−5√7+2√7−4√7 <------ who can solve this please!

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

thought you had this already..

OpenStudy (anonymous):

it is still \[3\sqrt{2}-7\sqrt{7}\]

OpenStudy (anonymous):

nooo one sayd,,that was the answer nd no one explained anyything nvm dang

OpenStudy (anonymous):

did you ask for an explanation <-- that's how u spell it rite..

OpenStudy (anonymous):

combine like terms: \[3x-5y+2y-4y\] \[3x+(-5+2-4)y\] \[3x-7y\]

OpenStudy (anonymous):

all the work is in \[-5+2-4=-7\]

OpenStudy (anonymous):

im sorry that was my brother i wen't to go outside and he got on my laptop he's in high school please forgive me for his actions

OpenStudy (anonymous):

um.. no problem..? i guess..? :s sorry for mine too.. woah im confused..

OpenStudy (anonymous):

lol,,it was my brother im sorry ,,now i knw i need to lock my laptop

OpenStudy (anonymous):

haha lool.. good luck!

OpenStudy (anonymous):

thanks lol,,he's a freshmen so he is still imature smh

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

haha lol gudnite!

OpenStudy (anonymous):

loool! nite!! dont let the bed bugs bite!